OpenStudy (anonymous):

Suppose that a vegetable vendor pays 7% interest daily on a $10 loan. Then, suppose that this individual was able to save $0.50 per day by drinking less tea. Calculate the number of days it would take for the vendor to no longer need to borrow money in order to purchase $10 of vegetables (i.e. the day number on which no loan is needed).

OpenStudy (anonymous):

n=the day number on which no loan is needed? r= interest daily=7%=0.07 Pv=(daily)=$0.5 Future Value=Fv=$10 n=ln(Fv/Pv)/ln(1+r) Using http://www.mathsisfun.com/money/compound-interest-calculator.html n=ln(10/0.5)/ln(1+0.07/)=

OpenStudy (anonymous):

2.995732274

OpenStudy (anonymous):

0.067658648

OpenStudy (anonymous):

29.49707588

OpenStudy (anonymous):

so i am guessing 29 days

OpenStudy (anonymous):

but which turns out to be wrong

OpenStudy (anonymous):

n= (log FV- log PV)/ log(1+r)